When applying for a professional position in a selected field, suppose that the probability of securing an interview is 0.15. A job seeker applies at 25 jobs in this field. What is the probability that 5 or more applications will result in interviews?
Answer #1

Let X be the random variable denoting the number of applications result in interviews out of 25 jobs.

The probability of securing an interview = 0.15

Thus, X ~ Bin(25, 0.15).

The probability that 5 or more applications will result in interviews out of 25 jobs = P(X \geq 5) = Ε 25 (0.15)(1 0.15) 25-1 Τ r-5 = 0.3179 (Ans).